OpenStudy (anonymous):

A stone falls freely from rest. How fast is it falling after 10 seconds?

OpenStudy (anonymous):

the stone starts frm rest so its initial velocity will be 0 as a =vf-vi/t a=vf/t as its falling so its speed will be constent at 9.8 m/ssquare a =9.8/10 a = .98m

OpenStudy (irishboy123):

use \[v = u + at, \ but \ u = 0, \ v = at, \ a = g = 9/8 \frac{m}{s^2}\] t = 10